# Newton-Flow Particle Filters based on Generalized Cramér Distance

### v1

@@ Abstract
Index Terms:
State estimation, particle filter, particle flow, homotopy continuation, progressive Bayes, differentiable particle filter, set distance, generalized Cramér distance, Newton flow.
@@
This is a preliminary version of the manuscript prepared as a basis for discussions during my plenary talk at the 2025 IEEE International Conference on Multisensor Fusion and Integration (MFI 2025) at Texas A&M University, College Station, Texas on September 3, 2025. The next version will include sections on implementation details, state-of-the-art, and an evaluation of the proposed filtering method.
@@ I-A Estimation Problem to be Solved
We consider recursive estimation of the state of a nonlinear dynamic system based on measurements sequentially arriving at discrete time steps. The evolution of the system state is described by a process model either in continuous time by a stochastic differential equation or in discrete time by a stochastic difference equation, both usually in state-space form. The relation of states to measurements is described by a stochastic measurement model. The goal is to continually provide an estimate of the true system state with each new measurement.
@@ I-C New Filtering Method: Key Ideas
The flow induced by the homotopy continuation very rarely produces exact results. In general, we have to be content with approximations. This means that we have to continually track the closest approximation to the true posterior over the artificial time. Closeness is measured by a set distance measure, the so called generalized Cramér-von Mises distance, that allows the efficient comparison of DMDs and is differentiable w.r.t. to the particle locations and weights. The optimal particle approximation is tracked over artificial time $\gamma$ by a Newton flow indced by this generalized Cramér-von Mises distance.
@@ I-D New Filtering Method: Unique Features
Here, we will give a short overview of the advantages of the new filtering method and also of its differences w.r.t the state-of-the-art.
@@
The new filtering method is different from other particle flow filter in that it does not calculate an explicit mapping for transporting particles. This can be interpreted as a direct online optimization of the distance measure versus finding a mapping policy first. This policy is often found by interpreting the homotopic version of the measurement update equation as a Fokker-Planck equation and using the associated (stochastic) ordinary differential equation (SDE) for moving the particles. This necessitates first finding the SDE and second numerically solving it for calculating the desired particle locations, which typically leads to numerical issues due to, e.g., stiffness. In contrast, the new method directly yields a closed-form ODE by calculating the gradient and Hessian of the generalized Cramér-von Mises distance. This ODE can be solved with standard methods.
@@
Methods relying on explicit mappings such as between the involved random variables are ultimately restricted to calculating posterior particle sets. The reason is that calculating the flow for Gaussian mixture densities (GMDs) or exponential densities based on the SDE for the posterior random variable (associated to the Fokker-Planck equation) would add another layer of complication. In fact, the SDE is a dynamical system describing the evolution of the posterior random variable and thus, not even weight changes of a particle set are possible (as the system mapping only changes sample locations, but not sample weights). In contrast, the new filtering method can be applied with different density representations. See [10] for an early approach involving GMDs.
@@ II-A Models
The models (1) and (2) relate random variables and their realization, and we call them generative models. Corresponding probabilistic models can be derived that describe the system evolution and the measurement extraction via conditional densities $f_{k}(\underline{x}_{k+1}|\underline{x}_{k})$ and $f_{k}(\underline{y}_{k}|\underline{x}_{k})$. For a specific measurement $\hat{\underline{y}}_{k}$, the so-called likelihood function $f_{k}^{L}$ is obtained by
$f_{k}^{L}(\underline{x})=f_{k}(\hat{\underline{y}}_{k}|\underline{x}_{k})\enspace.$ (3)
@@ II-C Estimation and Challenges
Please note that in this paper, we focus on a single filter step, omit the time index $k$, and change $f_{k}^{p}(\cdot)$, $f_{k}^{e}(\cdot)$ to $f_{p}(\cdot)$, $f_{e}(\cdot)$.
@@ II-E Challenges for General Density Representations
The filter step in (4) is exact and assumes that the posterior density can completely capture the shape of the true distribution. This works well for simple linear systems. For nonlinear systems, this only works when the posterior $\tilde{f}_{e}(\underline{x},\gamma)$ lives in a general infinite-dimensional distribution function space. For densities with limited representation capabilities, i.e., finite mixtures or finite particle sets, as required in practical implementations, the density type changes and/or the complexity goes up. For example, a GMD becomes a non-Gaussian mixture during the update or an equally weighted particle set becomes an unequally weighted particle set.
In recursive Bayesian estimation, we desire the densities describing the considered latent variables to be closed under the measurement update step, at least in an approximate sense. This means that for a prior density of a certain type, we want the posterior density to be of the same type [^1] Prior and posterior are then called (approximately) conjugate distributions.. In that case, we can use the resulting posterior as a prior for the next recursion step without changing its mechanism. This means that we have to find the best approximation of the resulting posterior lying within the space of permissible densities. This is equivalent to projecting the true posterior to the closest permissible approximation, which requires a distance measure.
@@ II-F Filter Step for DMDs
In this paper, we focus on the case that the prior density $f_{p}(\underline{x})$ is not a continuous density, but given as a set of $L$ weighted samples (or particles) it can be formally written as a DMD
$f_{p}(\underline{x})=\sum_{i=1}^{L}w_{p,i}\cdot\delta(\underline{x}-\underline{x}_{p,i})\enspace,$ (6)
@@
For prior DMDs, the posterior is also a DMD, and we can omit the normalization constant from the Bayesian filter step
$f_{e}(\underline{x})\propto f_{p}(\underline{x})\cdot f_{L}(\underline{x})\enspace,$ (7)
as normalizing $f_{e}(\underline{x})$ becomes trivial. Plugging in $f_{p}(\underline{x})$ from (6) gives
@@
$=\sum_{i=1}^{L}\underbrace{w_{p,i}\cdot f_{L}(\underline{x}_{p,i})}_{\bar{w}_{e,i}}\cdot\,\delta(\underline{x}-\underline{x}_{p,i})\enspace.$
@@
$\tilde{w}_{e,i}=\frac{\bar{w}_{e,i}}{\sum_{i=1}^{L}\bar{w}_{e,i}}\enspace.$ (9)
@@
$\tilde{f}_{e}(\underline{x})=\sum_{i=1}^{L}\tilde{w}_{e,i}\cdot\delta(\underline{x}-\tilde{\underline{x}}_{e,i})\enspace.$ (10)
@@ III Homotopy Continuation
In order to circumvent the problems coming with the one-shot measurement update shown in the previous section, [1] proposed to gradually introduce the information from the measurement. According to [2], this is done by carrying out the measurement update progressively over an artificial time $\gamma$, where $\gamma$ ranges in the (arbitrarily selected) interval $\gamma\in[0,1]$ [^2] Other intervals could be used when more convenient. This also includes $\gamma\in[0,\infty]$..
For achieving the desired effect, we parametrize the likelihood with the artificial time $\gamma$, resulting in $f_{L}(\underline{x},\gamma)$. For the initial value of $\gamma$, $\gamma=0$, a likelihood is selected that either allows a simple measurement update or that simply returns the prior. The latter case is equivalent to $f_{L}(\underline{x},\gamma)=c$, where we can select the constant $c$ to be $1$. For the final value of $\gamma$, $\gamma=1$, we want to recover the original likelihood function. In between the extreme values for $\gamma$, i.e., $\gamma\in(0,1)$, we desire an interpolation between the constant likelihood and the original likelihood that is convenient to use and leads to the desired estimation results.
@@
$f_{L}(\underline{x},\gamma)=\begin{cases}1&\gamma=0\\ \text{convenient interpolation}&\gamma\in(0,1)\\ f_{L}(\underline{x})&\gamma=1\end{cases}\enspace.$ (11)
Plugging the parametrized likelihood into the Bayesian update equation yields the true parametrized posterior
$\tilde{f}_{e}(\underline{x},\gamma)\propto f_{L}(\underline{x},\gamma)\,f_{p}(\underline{x})\enspace.$ (12)
There are infinitely many progression schedules and, strictly speaking, a good schedule used depends on the specific filtering setup, i.e., the prior density and the original likelihood. An intuitive and generic progression schedule is attained by taking the power of the original likelihood as
$f_{L}(\underline{x},\gamma)=\left[f_{L}(\underline{x})\right]^{\gamma}\enspace.$ (13)
@@
$f_{L}(\underline{x},\gamma)=\left[f_{L}(\underline{x})\right]^{g(\gamma)}\enspace.$ (14)
@@
$\dot{f}_{L}(\underline{x},\gamma)$ $=\frac{\partial}{\partial\,\gamma}\left\{\left[\exp\Big{(}\log\big{(}f_{L}(\underline{x})\big{)}\Big{)}\right]^{g(\gamma)}\right\}$ (15)
$=\frac{\partial}{\partial\,\gamma}\left\{\exp\Big{(}g(\gamma)\,\log\big{(}f_{L}(\underline{x})\big{)}\Big{)}\right\}$
$=\log\big{(}f_{L}(\underline{x})\big{)}\,\dot{g}(\gamma)\,f_{L}(\underline{x})\enspace.$
In many applications, we assume additive Gaussian measurement noise, which leads to likelihoods from the exponential family. For these likelihoods, we obtain simpler expressions due to the logarithm taken above. For additive zero-mean Gaussian noise with covariance matrix ${\mathbf{C}}_{v}$ and $g(\gamma)=\gamma$, the likelihood function is
$f_{L}(\underline{x},\gamma)=\exp\!\left(-\,\gamma\,\frac{1}{2}\big{(}\underline{y}-\underline{h}(\underline{x})\big{)}^{\top}{\mathbf{C}}_{v}^{-1}\big{(}\underline{y}-\underline{h}(\underline{x})\big{)}\right)\enspace.$ (16)
@@ IV Newton Flow Induced by Distance Measure
In a progressive Bayesian particle flow, a given intermediate posterior density undergoes continual changes for increasing artificial time $\gamma$ under the influence of the homotopic changes of the likelihood function. Hence, we require a distance measure that matches a slightly (infinitesimally) changed posterior density with an approximation from the desired density class. For particle sets, infinitesimal weight changes have to be instantly compensated by suitable location changes as we want to keep all weights equal.
@@ IV-A Abstract Distance Measure
We are given a reference density $\tilde{f}_{e}(\underline{x},\gamma)$ depending on the artificial time $\gamma$ and its approximation $f_{e}(\underline{x},\underline{\eta}(\gamma))$. The approximate posterior depends on a parameter vector $\underline{\eta}(\gamma)$ that implicitly depends on $\gamma$ as it changes with changes in $\gamma$.
An abstract distance measure for comparing $\tilde{f}_{e}(\underline{x},\gamma)$ and $f_{e}(\underline{x},\underline{\eta}(\gamma))$ is written as
$D\big{(}\underline{\eta}(\gamma),\gamma\big{)}=D\left(\tilde{f}_{e}(\underline{x},\gamma),f_{e}\left(\underline{x},\underline{\eta}(\gamma)\right)\right)\enspace.$ (17)
It is a function of $\underline{\eta}(\gamma)$ and $\gamma$, with an implicit dependency of $\underline{\eta}(\gamma)$ on $\gamma$.
As finding the optimal posterior is generally difficult, we exploit the homotopy from the previous section. Starting at the known minimum for $\gamma=0$, i.e., $\tilde{f}_{e}(\underline{x},\gamma=0)=f_{p}(\underline{x})$ and $f_{e}(\underline{x},\gamma=0)=\tilde{f}_{e}(\underline{x},\gamma=0)$, the minimum is tracked along the artificial time $\gamma$.
@@ IV-B Derivation of Iterative ODE
$\underline{G}\big{(}\underline{\eta}(\gamma),\gamma\big{)}=\frac{\partial D(\underline{\eta}(\gamma),\gamma)}{\partial\underline{\eta}(\gamma)}=\underline{0}\enspace.$ (18)
@@
$\underbrace{\frac{\partial\underline{G}(\underline{\eta}(\gamma),\gamma)}{\partial\underline{\eta}^{T}(\gamma)}}_{\underline{G}_{\underline{\eta}}\big{(}\underline{\eta}(\gamma),\gamma\big{)}}\cdot\frac{\partial\underline{\eta}(\gamma)}{\partial\gamma}+\underbrace{\frac{\partial\underline{G}\big{(}\underline{\eta}(\gamma),\gamma\big{)}}{\partial\gamma}}_{\underline{G}_{\gamma}(\underline{\eta}(\gamma),\gamma)}=\underline{0}\enspace$ (19)
@@
$\underbrace{\frac{\partial^{2}D\big{(}\underline{\eta}(\gamma),\gamma\big{)}}{\partial\underline{\eta}(\gamma)\,\partial\underline{\eta}^{T}(\gamma)}}_{D_{\underline{\eta}\underline{\eta}}(\underline{\eta}(\gamma),\gamma)}\cdot\underbrace{\frac{\partial\underline{\eta}(\gamma)}{\partial\gamma}}_{\dot{\underline{\eta}}(\gamma)}+\underbrace{\frac{\partial^{2}D\big{(}\underline{\eta}(\gamma),\gamma\big{)}}{\partial\underline{\eta}(\gamma)\,\partial\gamma}}_{D_{\underline{\eta}\gamma}(\underline{\eta}(\gamma),\gamma)}=\underline{0}\enspace.$ (20)
@@
$\underline{J}\big{(}\underline{\eta}(\gamma),\gamma\big{)}=D_{\underline{\eta}\gamma}\big{(}\underline{\eta}(\gamma),\gamma\big{)}=\underline{G}_{\gamma}\big{(}\underline{\eta}(\gamma),\gamma\big{)}$ (21)
@@
${\mathbf{H}}\big{(}\underline{\eta}(\gamma),\gamma\big{)}=D_{\underline{\eta}\underline{\eta}}(\underline{\eta}(\gamma),\gamma)=\underline{G}_{\underline{\eta}}(\underline{\eta}(\gamma),\gamma)\enspace.$ (22)
@@
${\mathbf{H}}\big{(}\underline{\eta}(\gamma),\gamma\big{)}\cdot\dot{\underline{\eta}}(\gamma)+\underline{J}\big{(}\underline{\eta}(\gamma),\gamma\big{)}=\underline{0}\enspace.$ (23)
It is an implicit ODE, where we have to solve for $\dot{\underline{\eta}}(\gamma)$. It can be formally rewritten as an explicit ODE
$\dot{\underline{\eta}}(\gamma)=-{\mathbf{H}}^{-1}\big{(}\underline{\eta}(\gamma),\gamma\big{)}\cdot\underline{J}\big{(}\underline{\eta}(\gamma),\gamma\big{)}\overset{\mathrm{def}}{=\joinrel=}\underline{R}\big{(}\underline{\eta}(\gamma),\gamma\big{)}\enspace.$ (24)
Please note that we desist from explicitly inverting the Hessian to convert the ODE into an explicit one, as numerical inversion has a high computational complexity and undesirable numerical properties. We rather solve for $\dot{\underline{\eta}}(\gamma)$ in (23) and exploit the symmetry of the Hessian.
We call this a Hessian flow for the parameter vector $\underline{\eta}(\gamma)$, in analogy to a gradient flow.
We also call this an iterative flow, as the reference posterior $\tilde{f}_{e}(\underline{x},\gamma)$ at time $\gamma$ is maintained as the product of the prior particle set times the progressive likelihood $f_{L}(\underline{x},\gamma)$. As the product of prior density times likelihood typically is not in the same density class as the prior, the reference posterior strays into the territory of impermissible densities and creeps beyond the border of usefulness, for significant changes in $\gamma$. To make this fact more visible, we can write the right-hand side $\underline{R}(\cdot,\cdot)$ in (24) as an explicit function of the reference posterior $\tilde{f}_{e}(\underline{x},\gamma)$ (and its derivative w.r.t. $\gamma$) as
$\dot{\underline{\eta}}(\gamma)=\underline{R}\big{(}\underline{\eta}(\gamma),\tilde{f}_{e}(\underline{x},\gamma),\dot{\tilde{f}}_{e}(\underline{x},\gamma)\big{)}\enspace.$ (25)
In order to avoid the reference posterior becoming useless, we have to replace it with its approximation either from time to time or continually. We will call the resulting ODE a recursive flow.

### v2

@@ Abstract
The first version of this manuscript was prepared as a basis for discussions during my plenary talk at the 2025 IEEE International Conference on Multisensor Fusion and Integration (MFI 2025) at Texas A&M University, College Station, Texas on September 3, 2025.
This slightly modified version was prepared for my plenary talk at the 17th Symposium Sensor Data Fusion in Bonn, Germany, on November 25, 2025. It includes changes based on comments by Ryne Beeson, Ondřej Straka, Daniel Frisch, and Yiqing Wang. In addition, the distance measure was modified to naturally avoid singularities.
The next version will include sections on implementation details, state-of-the-art, and an evaluation of the proposed filtering method.
@@ I-A Estimation Problem to be Solved
We consider recursive estimation of the state of a nonlinear dynamic system based on measurements sequentially arriving at discrete time steps. The evolution of the system state is described by a process model either in continuous time by a stochastic differential equation (SDE) or in discrete time by a stochastic difference equation, both usually in state-space form. The relation of states to measurements is described by a stochastic measurement model. The goal is to continually provide an estimate of the true system state with each new measurement.
@@ I-C New Filtering Method: Key Ideas
The flow induced by the homotopy continuation very rarely produces exact results. In general, we have to be content with approximations. This means that we have to continually track the closest approximation to the true posterior over the artificial time. Closeness is measured by a set distance measure, the so called generalized Cramér-von Mises distance [7, 10] or Cramér distance for short, that allows the efficient comparison of DMDs and is differentiable w.r.t. to the particle locations and weights. The optimal particle approximation is tracked over artificial time $\gamma$ by a Newton flow induced by this generalized Cramér distance.
@@ I-D New Filtering Method: Unique Features
Here, we will give a short overview of the advantages of the new filtering method and also of its differences w.r.t. the state-of-the-art.
@@
The new filtering method is different from other particle flow filters in that it does not calculate an explicit mapping for transporting particles. This can be interpreted as a direct online optimization of the distance measure versus finding a mapping policy first. Such a policy is often found by interpreting the homotopic version of the measurement update equation as a Fokker-Planck equation and using the associated SDE for moving the particles. This necessitates first finding the SDE and second numerically solving it for calculating the desired particle locations, which typically leads to numerical issues due to various reasons, e.g., stiffness. In contrast, the new method directly yields a closed-form ODE by calculating the gradient and Hessian of the generalized Cramér distance. This ODE can be solved with standard methods.
@@
Methods relying on explicit mappings between the involved random variables are ultimately restricted to calculating posterior particle sets. The reason is that calculating the flow for Gaussian mixture densities (GMDs) or exponential densities based on the SDE for the posterior random variable (associated to the Fokker-Planck equation) would add another layer of complication. In fact, the SDE is a dynamic system describing the evolution of the posterior random variable and thus, not even weight changes of a particle set are possible (as the system mapping only changes sample locations, but not sample weights). In contrast, the new filtering method can be applied with different density representations. See [11] for an early approach involving GMDs.
@@ II-A Models
The models (1) and (2) relate random variables and their realizations, and we call them generative models. Corresponding probabilistic models can be derived that describe the system evolution and the measurement extraction via conditional densities $f_{k}(\underline{x}_{k+1}|\underline{x}_{k})$ and $f_{k}(\underline{y}_{k}|\underline{x}_{k})$. For a specific measurement $\hat{\underline{y}}_{k}$, the so-called likelihood function $f_{k}^{L}$ is obtained by
$f_{k}^{L}(\underline{x})=f_{k}(\hat{\underline{y}}_{k}|\underline{x}_{k})\kern 5.0pt.$ (3)
@@ II-C Estimation and Challenges
Please note that in this paper, we focus on a single filter step, omit the time index $k$, and change $f_{k}^{p}(\cdot)$, $f_{k}^{e}(\cdot)$, $f_{k}^{L}(\cdot)$ to $f_{p}(\cdot)$, $f_{e}(\cdot)$, $f_{L}(\cdot)$.
@@ II-E Challenges for General Density Representations
The filter step in (4) is exact and assumes that the posterior density can completely capture the shape of the true distribution. This is true, e.g., in linear systems. For nonlinear systems, it only works when the posterior $f_{e}(\underline{x})$ lives in a general infinite-dimensional distribution function space. For densities with limited representation capabilities, i.e., finite mixtures or finite particle sets, as required in practical implementations, the density type changes and/or the complexity goes up. For example, a GMD becomes a non-Gaussian mixture during the update or an equally weighted particle set becomes an unequally weighted particle set.
In recursive Bayesian estimation, we desire the densities describing the considered latent variables to be closed under the measurement update step, at least in an approximate sense. This means that for a prior density of a certain type, we want the posterior density to be of the same type [^1] Prior and posterior are then called (approximately) conjugate distributions w.r.t. a likelihood.. In that case, we can use the resulting posterior as a prior for the next recursion step without changing its mechanism. This means that we have to find the best approximation of the resulting posterior lying within the space of permissible densities. This is equivalent to projecting the true posterior to the closest permissible approximation, which requires a distance measure.
@@ II-F Filter Step for DMDs
In this paper, we focus on the case that the prior density $f_{p}(\underline{x})$ is not a continuous density, but given as a set of $L$ weighted samples (or particles). It can be formally written as a DMD
$f_{p}(\underline{x})=\sum_{i=1}^{L}w_{p,i}\cdot\delta(\underline{x}-\underline{x}_{p,i})\kern 5.0pt,$ (6)
@@
For prior DMDs, the posterior is also a DMD, and we can omit the normalization constant from the Bayesian filter step (4)
$\tilde{f}_{e}(\underline{x})\propto f_{L}(\underline{x})\cdot f_{p}(\underline{x})\kern 5.0pt,$ (7)
as normalizing $\tilde{f}_{e}(\underline{x})$ becomes trivial. The tilde $\tilde{\phantom{\ast}}$ marks this as the true posterior that will be approximated later. Plugging in $f_{p}(\underline{x})$ from (6) gives
@@
$=\sum_{i=1}^{L}\underbrace{w_{p,i}\cdot f_{L}(\underline{x}_{p,i})}_{\bar{w}_{e,i}}\cdot\,\delta(\underline{x}-\underline{x}_{p,i})\kern 5.0pt.$
@@
$\tilde{w}_{e,i}=\frac{\bar{w}_{e,i}}{\sum_{i=1}^{L}\bar{w}_{e,i}}\kern 5.0pt.$ (9)
@@
$\tilde{f}_{e}(\underline{x})=\sum_{i=1}^{L}\tilde{w}_{e,i}\cdot\delta(\underline{x}-\tilde{\underline{x}}_{e,i})\kern 5.0pt.$ (10)
@@ III Homotopy Continuation
In order to circumvent the problems coming with the one-shot measurement update shown in the previous section, [1] proposed to gradually introduce the information from the measurement. According to [2], this is done by carrying out the measurement update progressively over an artificial time $\gamma$, where $\gamma$ ranges in the (arbitrarily selected) interval [^2] Other intervals could be used when more convenient. This also includes $\gamma\in[0,\infty]$. $\gamma\in[0,1]$.
To achieve the desired effect, we parametrize the likelihood with the artificial time $\gamma$, resulting in $f_{L}(\underline{x},\gamma)$. For the initial value of $\gamma$, $\gamma=0$, an uninformative likelihood, e.g., $f_{L}(\underline{x},\gamma=0)=1$, is selected that simply returns the prior unchanged. For the final value of $\gamma$, $\gamma=1$, we want to recover the original likelihood function. In between the extreme values for $\gamma$, i.e., $\gamma\in(0,1)$, we desire an interpolation between the constant likelihood and the original likelihood that is convenient to use and leads to the desired estimation results.
@@
$f_{L}(\underline{x},\gamma)=\begin{cases}1&\gamma=0\\ \text{convenient interpolation}&\gamma\in(0,1)\\ f_{L}(\underline{x})&\gamma=1\end{cases}\kern 5.0pt.$ (11)
Plugging the parametrized likelihood into the Bayesian update equation (7) yields the true parametrized posterior
$\tilde{f}_{e}(\underline{x},\gamma)\propto f_{L}(\underline{x},\gamma)\,f_{p}(\underline{x})\kern 5.0pt.$ (12)
There are infinitely many progression schedules and, strictly speaking, a good schedule depends on the specific filtering setup, i.e., the prior density and the original likelihood. An intuitive and generic progression schedule is attained by taking the power of the original likelihood as
$f_{L}(\underline{x},\gamma)=\left[f_{L}(\underline{x})\right]^{\gamma}\kern 5.0pt.$ (13)
@@
$f_{L}(\underline{x},\gamma)=\left[f_{L}(\underline{x})\right]^{g(\gamma)}\kern 5.0pt.$ (14)
@@
$\dot{f}_{L}(\underline{x},\gamma)$ $=\frac{\partial}{\partial\,\gamma}\left\{\left[\exp\Bigl(\log\bigl(f_{L}(\underline{x})\bigr)\Bigr)\right]^{g(\gamma)}\right\}$ (15)
$=\frac{\partial}{\partial\,\gamma}\left\{\exp\Bigl(g(\gamma)\,\log\bigl(f_{L}(\underline{x})\bigr)\Bigr)\right\}$
$=\log\bigl(f_{L}(\underline{x})\bigr)\,\dot{g}(\gamma)\,f_{L}(\underline{x},\gamma)\kern 5.0pt.$
In many applications, we assume additive Gaussian measurement noise, which leads to likelihoods from the exponential family. For these likelihoods, we obtain simpler expressions due to the logarithm taken above. For additive zero-mean Gaussian noise with covariance matrix ${\mathbf{C}}_{v}$ and $g(\gamma)=\gamma$, the likelihood function corresponding to (2) is
$f_{L}(\underline{x},\gamma)=\exp\!\left(-\,\gamma\,\frac{1}{2}\bigl(\underline{y}-\underline{h}(\underline{x})\bigr)^{\top}{\mathbf{C}}_{v}^{-1}\bigl(\underline{y}-\underline{h}(\underline{x})\bigr)\right)\kern 5.0pt.$ (16)
@@ IV Newton Flow Induced by Distance Measure
In a progressive Bayesian particle flow, a given intermediate posterior density undergoes continual changes for increasing artificial time $\gamma$ under the influence of the homotopic changes of the likelihood function. Hence, we require a distance measure that matches a slightly (infinitesimally) changed posterior density with an approximation from the desired density class. For particle sets, infinitesimal weight changes have to be instantly compensated by suitable particle location changes as we want to keep all weights equal.
@@ IV-A Abstract Distance Measure
We are given a true or reference posterior density $\tilde{f}_{e}(\underline{x},\gamma)$ explicitly depending on the artificial time $\gamma$ and an approximation $f_{e}(\underline{x},\underline{\eta}(\gamma))$ of the reference density. The parameter vector $\underline{\eta}(\gamma)$ is adjusted over time $\gamma$ so that the approximate posterior follows the evolving reference density. As a result, $\underline{\eta}(\gamma)$ implicitly depends on $\gamma$.
To calculate the necessary changes in the parameter vector $\underline{\eta}(\gamma)$, an abstract distance measure for comparing $\tilde{f}_{e}(\underline{x},\gamma)$ and $f_{e}(\underline{x},\underline{\eta}(\gamma))$ is used
$D\bigl(\underline{\eta}(\gamma),\gamma\bigr)=D\left(\tilde{f}_{e}(\underline{x},\gamma),f_{e}\left(\underline{x},\underline{\eta}(\gamma)\right)\right)\kern 5.0pt.$ (17)
Changing $\gamma$ induces variations in the distance that have to be compensated by changes in $\underline{\eta}(\gamma)$ in order to stay in the minimum.
Finding the optimal approximate posterior density by minimizing the distance measure (17) is generally difficult due to local minima. Hence, we start at the known minimum for $\gamma=0$, i.e., $f_{e}(\underline{x},\gamma=0)=\tilde{f}_{e}(\underline{x},\gamma=0)=f_{p}(\underline{x})$, and track the optimal $\gamma$-dependent approximate posterior $f_{e}(\underline{x},\underline{\eta}(\gamma))$ as the minimum of (17) with infinitesimal changes of the artificial time $\gamma$. The desired final approximate posterior is obtained at $\gamma=1$.
@@ IV-B Derivation of Iterative ODE
$\underline{G}\bigl(\underline{\eta}(\gamma),\gamma\bigr)=\underbrace{\frac{\partial D(\underline{\eta}(\gamma),\gamma)}{\partial\underline{\eta}(\gamma)}}_{D_{\underline{\eta}}(\underline{\eta}(\gamma),\gamma)}=\underline{0}\kern 5.0pt.$ (18)
@@
$\underbrace{\frac{\partial\underline{G}(\underline{\eta}(\gamma),\gamma)}{\partial\underline{\eta}^{T}(\gamma)}}_{\underline{G}_{\underline{\eta}}\bigl(\underline{\eta}(\gamma),\gamma\bigr)}\cdot\frac{\partial\underline{\eta}(\gamma)}{\partial\gamma}+\underbrace{\frac{\partial\underline{G}\bigl(\underline{\eta}(\gamma),\gamma\bigr)}{\partial\gamma}}_{\underline{G}_{\gamma}(\underline{\eta}(\gamma),\gamma)}=\underline{0}\kern 5.0pt$ (19)
@@
$\underbrace{\frac{\partial^{2}D\bigl(\underline{\eta}(\gamma),\gamma\bigr)}{\partial\underline{\eta}(\gamma)\,\partial\underline{\eta}^{T}(\gamma)}}_{D_{\underline{\eta}\underline{\eta}}(\underline{\eta}(\gamma),\gamma)}\cdot\underbrace{\frac{\partial\underline{\eta}(\gamma)}{\partial\gamma}}_{\dot{\underline{\eta}}(\gamma)}+\underbrace{\frac{\partial^{2}D\bigl(\underline{\eta}(\gamma),\gamma\bigr)}{\partial\underline{\eta}(\gamma)\,\partial\gamma}}_{D_{\underline{\eta}\gamma}(\underline{\eta}(\gamma),\gamma)}=\underline{0}\kern 5.0pt.$ (20)
@@
$\underline{J}\bigl(\underline{\eta}(\gamma),\gamma\bigr)=D_{\underline{\eta}\gamma}\bigl(\underline{\eta}(\gamma),\gamma\bigr)=\underline{G}_{\gamma}\bigl(\underline{\eta}(\gamma),\gamma\bigr)$ (21)
@@
${\mathbf{H}}\bigl(\underline{\eta}(\gamma),\gamma\bigr)=D_{\underline{\eta}\underline{\eta}}(\underline{\eta}(\gamma),\gamma)=\underline{G}_{\underline{\eta}}(\underline{\eta}(\gamma),\gamma)\kern 5.0pt.$ (22)
@@
${\mathbf{H}}\bigl(\underline{\eta}(\gamma),\gamma\bigr)\cdot\dot{\underline{\eta}}(\gamma)+\underline{J}\bigl(\underline{\eta}(\gamma),\gamma\bigr)=\underline{0}\kern 5.0pt.$ (23)
It is an implicit ODE with mass matrix ${\mathbf{H}}(\cdot,\cdot)$, where we have to solve for $\dot{\underline{\eta}}(\gamma)$. It can be formally rewritten as an explicit ODE
$\dot{\underline{\eta}}(\gamma)=-{\mathbf{H}}^{-1}\bigl(\underline{\eta}(\gamma),\gamma\bigr)\cdot\underline{J}\bigl(\underline{\eta}(\gamma),\gamma\bigr)\overset{\mathrm{def}}{=\joinrel=}\underline{R}\bigl(\underline{\eta}(\gamma),\gamma\bigr)\kern 5.0pt.$ (24)
Please note that we desist from explicitly inverting the Hessian to convert the ODE into an explicit one, as numerical inversion has a high computational complexity and undesirable numerical properties. We rather solve for $\dot{\underline{\eta}}(\gamma)$ in the linear system (23) while exploiting the properties of the Hessian.
We call this a Newton flow for the parameter vector $\underline{\eta}(\gamma)$ as a continuous version of Newton’s method, in analogy to a gradient flow being a continuous version of a gradient-descent method.
We also call this an iterative flow, as the reference posterior $\tilde{f}_{e}(\underline{x},\gamma)$ at time $\gamma$ is maintained as the product of the prior particle set times the progressive likelihood $f_{L}(\underline{x},\gamma)$. As the product of prior density times likelihood typically is not in the same density class as the prior, the reference posterior $\tilde{f}_{e}(\underline{x},\gamma)$ strays into the territory of impermissible densities and, for significant changes in $\gamma$, creeps beyond the border of usefulness. To make the effect on the flow in (24) more visible, we note that changes in $\gamma$ enter the right-hand side $\underline{R}(\cdot,\cdot)$ via the reference posterior $\tilde{f}_{e}(\underline{x},\gamma)$ and its derivative $\dot{\tilde{f}}_{e}(\underline{x},\gamma)$ w.r.t. $\gamma$. This means that we can write $\underline{R}(\cdot,\cdot)$ as an explicit function of $\tilde{f}_{e}(\underline{x},\gamma)$ and $\dot{\tilde{f}}_{e}(\underline{x},\gamma)$ as
$\dot{\underline{\eta}}(\gamma)=\underline{R}\bigl(\underline{\eta}(\gamma),\tilde{f}_{e}(\underline{x},\gamma),\dot{\tilde{f}}_{e}(\underline{x},\gamma)\bigr)\kern 5.0pt.$ (25)
In order to avoid the reference posterior becoming useless, we have to replace it with its approximation $f_{e}(\underline{x},\underline{\eta}(\gamma))$ either from time to time or continually. We will call the resulting ODE a recursive flow.